Bookmark this page Setup: All problems on this page will follow the definitions here: Let X, Y be two Bernoulli random variabNote: In all problems below, do not assume r = pq. Define Ρ - ΣΧ, 4 = ΗΣΥAsymptotic Variance Under the Null 1.5 points possible (graded, results hidden) Consider the following hypotheses: H. : X and
Answer #1

Ang:- Guien P= p(x:1) q: p(y=1) Y : P(x-1, y-1) %3D Tn Calculalion :- let x,Y be ló Bernoil vrandon in and ib Y- Pg hen x and2. PCANB) = P(A) - P(ANB) NOW KA) = P, P(AnB) =} * PCAns)- P-8 - P(x=1,y=0) -> P(x=0)(Y=1) pado is P(AnB) Reg P(B) - P(BNA)3) NOw PCANB) = |- p( AUB) = | - KA)- P(B) + P(aNB) = |-p-9+Y P(x=0,Y-0) = |- p-9+Y %3D ib A and B îndepenoent eent Hence th